OpenStudy (anonymous):

Fill in the blank to make the expression a perfect square w^2-6w+

OpenStudy (amistre64):

(x+a)^2 is a perfect square (x+a)^2 = x^2 + 2ax + a^2 -6 = 2a a = -6/2 so what is a^2?